OpenStudy (anonymous):

Graph the equation with a diameter that has endpoints (-3, 4) and (5, -2). Label the center and at least four points on the circle. Write the equation of the circle. Help please?

OpenStudy (ddcamp):

Use the midpoint formula to find the center: \[(\frac{ x₂+x₁ }{ 2 } ,\frac{ y₂+y₁ }{ 2 })\] Use that center point and one of the endpoints to find the radius (distance formula) Then, plug it in to the formula for a circle: \[(x-h)²+(y-k)² = r²\] Where (h,k) is the center point and r is the radius.

OpenStudy (anonymous):

I got the coordinates (1,1) (x - 1)² + (y - 1)² = 5²

OpenStudy (ddcamp):

The center is at (1,1), and the radius is 5. So yes, that's correct!